# PizzaCommonSense: Learning to Model Commonsense Reasoning about Intermediate Steps in Cooking Recipes

### v1

@@ 2 Related Work
Large datasets focusing on the cooking domain have garnered new interest, contributing to advancements in food understanding. The most notable examples of such datasets are Food-101Bossard et al. (2014) and Recipe1M Salvador et al. (2017a). However, these benchmark datasets have mostly been exploited as reference benchmarks for computer vision, e.g. see Pham et al. (2021). The text-based studies in natural language processing have focused on tasks such as recipe text understanding in the form of flow-graphs Mori et al. (2014a, b) and more generally recipe parsing Chang et al. (2018); Jermsurawong and Habash (2015); Kiddon et al. (2015). We propose a dataset to explain the input-output pairs and state changes in the cooking process in natural language. The supervision provided by these fine-grained, step by step, annotations aims at obtaining models that are capable of reasoning and explaining their reasoning path. This aspect is of particular importance for the goal towards generating safe and coherent recipes.
Procedural text understanding
Procedural text understanding is a challenging subfield of natural language processing (NLP) that focuses on understanding the step-by-step processes described in natural language Bosselut et al. (2018); Henaff et al. (2019). This task involves extracting the entities, actions, and relationships involved in the process, as well as understanding the temporal and causal dependencies between them. Different approaches to tackle the task have been proposed. For example, syntactic parsing methods such as AMR (abstract meaning representation) Banarescu et al. (2013) represent “who did what to whom” by uncovering stated facts, or again tracking entity states uncovers unstated facts such as how ingredients change during a recipe. However, PizzaCommonSense goes further by explicitly predicting the input/output pairs of each action with fine-grained annotations. These input/output sequences are formulated in natural language and serve to explain the reasoning behind the actions. This type of supervision could enhance the clarity, understandability, and reusability of the text. By clearly defining the conditions under which each action is performed and the resulting state it produces, the text becomes more self-explanatory, reducing the need for external context or assumptions. This explicitness also makes the text more amenable to automation, since software agents can easily interpret and execute the actions described.
@@
into smaller, more manageable subtasks is an effective approach for training models to tackle complex problems. This decomposition allows the model to focus on specific aspects of the task and to gradually build an understanding of the overall problem. Most recent approaches to these problems are shifting towards using LLMs, often relying on approaches involving prompting or in-context learning. One such prompting approach is the chain of thought prompting Wei et al. (2022), which prompts the language model to generate a series of intermediate steps that improve the reasoning capabilities in LLMs. Wang et al. took another step forward and sampled multiple reasoning paths and selected the most relevant output using majority voting. Kojima et al. further improved the reasoning of LLMs in a zero-shot manner by appending “Let’s think step by step” to the prompt. In contrast, our work does not propose prompting solutions; instead, we explicitly supervise the learning model to reason by trying to solve the sub-question at each step. Most similar to our work is the work of Zhou et al. which decomposes questions into sub-questions and asks the language model to solve each sub-question sequentially or Shridhar et al. (2023) which proposes to distill reasoning capabilities into smaller models and suggests ways to further improve them by explicitly asking stepwise questions.
@@
The reasoning process consisting of decomposing a multi-step reasoning task is inherently structured. While conventional natural language texts are generated in a 1-dimensional sequential order, the table has a 2-dimensional structure, which allows to reason horizontally and vertically at the same time. We argue that these features justify the choice of tabular data for the proposed task. To use an LLM for tabular data, the table must be serialized into a natural text representation. Proposed serialization formats include the simple list or sentence serializations Narayan et al. (2022); Borisov et al. (2022). Yin et al. (2020) also included the column data type in the serialized string. We chose the serialization proposed by Wu et al. (2022).
3 Dataset
3.1 Data preparation
We build our dataset on Recipe1M by Salvador et al. (2017b) which consists of one million structured cooking recipes with 13M associated images. We sample a set of 1087 recipes that contain the word "pizza" in the title. It is worth mentioning that despite the choice of selecting the recipes explicitly mentioning the word pizza, not all recipes have as a final product a pizza. In order to adapt the recipes to our task, we first split each sentence in the recipes such that it contains only one relevant cooking action. We parse the sentences such that only coordinate sentences are split, for example, "Heat and stir to mix evenly" becomes two separate instructions "Heat" and "stir to mix heavenly": "heat" is the ROOT, "stir" is connected to the root by a coordinating conjunction, hence of same importance. On the other hand, "mix" is preceded by the particle "to" and acts as a clause modifier. In another example, the sentence "after you heat, stir to mix evenly" will not be split as "after you heat" is a prepositional phrase. The cooking actions glossary is obtained from the dataset Now You’re Cooking Kiddon et al. (2016); Bosselut et al. (2018). This choice is motivated by the goal of this dataset to promote reasoning about the purpose of an action and the intermediate transformations necessary to prepare a given dish. Hence input-output pairs are reasonable only for actions that promote a state change in the comestibles at hand.
Annotation collection
We ran the annotation process on Amazon Mechanical Turk (AMT). We choose to frame the data collection process in a tabular format which is well adapted to the task for the clarity and conciseness and makes it easier to understand the relationships between the information in the columns and the different rows. Organizing the information using this format allows easy identification and analysis of the relationships between input and output, as well as those between different instructions. This format makes the annotation collection more user-friendly for the crowd workers.
The columns of interest are: Instruction, Input, Action, and Output. We ask the crowd-workers to provide clear and understandable descriptions of the intermediates given the action and the instructions. We enumerate some constraints:
  • The "input" cell represents the state of the food preparation before the cooking action is performed and the "output" cell represents the state of the food preparation after the cooking action is performed.
  • For steps that do not refer to any comestibles (e.g. preheat the oven to 450F), "NA" need to be used for input and output. Comestibles in a set are separated by a semicolon. This occurs several times in the input column.
  • Verbs of motion such as move, place, transfer will have same input and output.
The first row of each table is pre-filled, if the cooking action is a verb whose object is not a comestible with "NA". In addition, the crowd-workers performing the HIT were asked to not insert numerical values and if the output of a previous step were to become the input of a following instruction, to keep the description the same.
@@ 3.1 Data preparation
Key statistics
Table 1 shows the key statistics of PizzaCommonsense. The dataset contains 13141 data instances of instruction, input, cooking action and output among 1087 annotated recipes. An annotated example is shown in figure 2. The average completion time per recipe is 5 minutes. We do not collect personal information about the crowd-workers. We manually checked the annotated recipes to ensure the quality of the collected data.
Distribution-based split
In order to ensure that the distribution of ingredients is preserved across the training, validation and test sets we use distribution-based data splits obtained through clustering. First, we extract the ingredients of each recipe, process them in order to ignore specific terms such as brands and quantities, and then vectorize them. The clustering algorithm iteratively merges recipes to the clusters based on a distance metric. Finally, we assign the recipes from each cluster to the respective splits, ensuring an accurate representation and evaluation.
4 Task
We extend the goal of procedural text comprehension with an additional objective of identifying the corresponding input intermediate state and output intermediate state for each instruction step. In other words, the goal is to develop a model that can effectively understand sequential instructions at a fine-grained level. The model should be able to accurately predict both the input and output for each instruction, including the implicit and explicit ingredients involved, the outcomes of the actions performed, and the resulting condition of each step. We argue that this ability will lead to the generation of faithful variations of the recipe and the ability to reason about the recipe’s structure and components.
Problem Formalization
We have a tabular representation of a recipe $R$ with $n=4$ columns (Instructions, Input, Action, Output), which also corresponds to the natural language strings representing the table header, and $m$ rows, one for each instruction of the recipe. The goal is to predict the content of the columns Input and Output, given the content of the columns Instruction and Action. An extra requirement of these tables are that (i) the output cell of the row $t-1$ can be the content of the input cell of the row $t$ and (ii) the input and output cells of a row $t$ can be the same, if the cooking action is not transformative.
Serialization of Tabular Data
The proposed baselines are based on language models, hence the tables at hand must be transformed into a natural text representation. We define a function serialize(R) that takes as input the tabular representation of the recipe $R$ and outputs a textual representation of the input. To handle the missing values in the table, the function mask fills the missing values with a mask token according to the architecture at hand. For example, for the baseline T5, we rely on the predefined sentinel tokens for the original span denoising objective. Otherwise, we use standard masking tokens <in> and <out> to indicate the missing values. The <s> token is used to separate the cells of each row, <n> is used to separate the rows in the serialized table. The content of the input and output cells can be a set of comestibles or ingredients. For example, the instruction explaining how to assemble a dish is likely to have multiple food items that need to be combined. In order to ensure clarity and readability, the ingredients are listed between brackets and separated by semicolons. An example of this is given in Fig.2.
Input Output
EMA R R BS
T5 0.2 0.71 0.84 59.9
Flan-T5 0.0 0.0 0.64 26.2
GPT-3.5 26.7 51.4 52.3 88.9
[…]
Random Clustering
Input Output Input Output
EMA R B R M BS EMA R B R M BS
[…]
4.1 Fine Tuning T5-based models
T5
Raffel et al. (2020) is a generative pre-trained transformer model that can be used for a wide variety of natural language understanding (NLU) tasks such as summarization, and question answering. During the span mask denoising objective, T5 is trained on a massive dataset of text that has been corrupted with spans of tokens removed. The goal of this objective is to teach T5 to reconstruct the original text from the corrupted input. To achieve this, T5 is trained to predict the sentinel tokens, which are special tokens that are inserted into the corrupted input to mark the positions where tokens have been removed. The predicted tokens are then placed between the sentinel tokens to restore the original text. We take advantage of this pretrained objective and we build a baseline through fine tuning the pretrained T5-base model in a sequence to sequence fashion. For doing this, we use as input the serialized table with the input and output masked out and the output are the content of corresponding input and output.
Flan T5
Additionally, we use Flan T5 as a baseline for our approach. It is an enhanced version of T5 that has been fine-tuned in a mixture of tasks using a technique called instruction-finetuning. To take advantage of this fact, we append the prompt "Fill in the placeholders <in> and <out> with the appropriate comestible or food preparation." to guide the fine-tuning process. The rest of the experiment is then conducted in the same way as the T5 based approach.
4.2 GPT-3.5
In-context learning
Due to the specificity of the structure of the task, we investigate in-context learning setting for predicting the input/output pairs. To do this, for each test sample (serialize(R), serialize(mask(R))) from the annotated dataset, we sample a recipe table from the training set which is used as demonstration for GPT-3.5 to perform the task. Contrary to the T5-based experiments, the output given is the completed serialized table.
Fine-tuning
Due to the specificity of the structure of the task, we investigate fine tuning GPT3.5. Fine-tuning is a crucial step in tailoring the language model’s capabilities to specific tasks and domains. By exposing the model to a targeted dataset of relevant examples, fine-tuning enhances its ability to generate accurate and relevant text that aligns with the desired context and purpose. We follow the same structure used for fine-tuning T5 where the input and output are masked with special tokens <in> and <out>.
4.3 Preliminary study
We first conduct a preliminary study to inspect the performance of the chosen baselines without fine-tuning. After serializing the table, we feed each table in our annotated test set to the baseline models. From Table 2 we can observe that all models perform poorly out-of-the-box. In particular, T5 models fail to predict appropriate text. On the other hand, GPT-3.5 failed to generate predictions in the correct tabular format in 72% of the cases. In those cases, the model will either paraphrase or rewrite the instructions of recipe, or predict the title and/or the final preparation object of the recipe, or again predict (part of) the correct placeholders in a different format. An example of this last case is shown in the appendix D. All those are ignored while computing the preliminary results.
5 Evaluation
5.1 Evaluation metrics
[…]
5.2 Experimental setting
We use T5-base and Flan T5-base models with a batch size of 16 and a learning rate of $1e^{-5}$, using Adam optimizer Kingma and Welling (2013). We fine-tune for 30 epochs and save the checkpoint with the best performance on the validation set. Both T5 based models are pretrained in English. Both T5 and Flan T5 employs around an hour on 4 GTX GPUs. For the LLM-based method, we use gpt-3.5-turbo-1106 model and evaluate it on the annotated test set. The cost of finetuning the model was 15 USD and the cost per API call for the other configurations was around 2 USD.
6 Results
We analyze the performance of the baseline models in predicting the correct input and output for the intermediate steps of a cooking recipe. To perform well, the models should predict (i) only comestibles in the input/output pairs; (ii) all the comestibles necessary to perform the cooking action at the given time step, implicitly and explicitly stated in the cooking instruction. In particular, it should not fail to include the output of the previous step if the next cooking instruction is a transformation of the comestible; (iii) the comestibles implicitly mentioned could also be conveyed by the cooking action, i.e. the input associated with Salt the sauce should be (salt; sauce) and finally (iv) specific descriptions of the output comestibles, i.e. although not strictly wrong, describing a sauce as a mixture of tomato paste, oil, water and seasonings is undesirable.
@@ 6 Results
+ demo 22.3 36.9 32.5 9.20 20.0 87.2
+ FT 32.6 55.9 53.6 15.8 46.2 90.6
[…]
6.1 T5 and Flan-T5
T5
The evaluation results of the T5-based methods are shown in Table 3. All the methods show imperfect results, which suggests that the proposed task is challenging. In particular, the random splits setting exhibits better scores although not by a large margin compared to the distribution based setting. Comparing the two settings, we can see that T5 doesn’t generalize well and suffers from a drop in performance of 21% in exact match accuracy for what concerns the input and 17% in Bleu scores for the output description. It’s worth adding that both the pretrained versions of T5 and Flan T5 failed to give any meaningful prediction. For this type of baseline, the fine-tuning option was a necessity. We can observe that Flan-T5 improves in almost all scores compared to T5. Although there still is a decrease in performance from the distribution based split to the random setting, the BertScore is consistently higher than the T5 counterpart. This suggests that fluency, semantic meaning and contextual coherence of the generated description is improved.
6.2 GPT-3.5
[…]
Instruction Ground-truth T5 Flan T5 GPT+demo GPT
(tomato sauce; water; oregano; …)
(chicken; oil; cheese) (all ingredient; oil; cheese) tomatoes, onion, garlic, oregano,… (tomato; salt; italian seasoning)
combine all ingredients except for the chicken , oil and cheese in a saucepan.
seasoned tomato sauce
sauce chicken fat combined tomato sauce mixture tomato sauce
(opt. flour; opt. water; dough)
(dough; flour; water) (mixed ingredients; flour; water) flour or water (dough; water or flour)
add more flour or water until you get that consistency
dough
flour added to dough flour or water added to mixed ingredients adjusted consistency water or flour added to dough
(…; salt; blue cornmeal)
(dough; olive oil) (peanut butter; fresh dill; …) ingredients ingredients
place everything in the bowl of an electric mixer with a dough hook
partially mixed blue dough
dough peanut butter mixture ingredients in the bowl ingredients
NA
(chicken; fork) NA NA (fork; chicken)
before putting the chicken on , get a fork
NA
chicken with fork NA NA chicken and fork
7 Analysis & Discussion
The evaluation of the baselines’ performance revealed both promising outcomes and areas for improvement. This discussion delves into the details of the results, identifying strengths and weaknesses in the model’s predictions. By understanding the factors that contribute to both accurate and erroneous predictions, we can gain valuable insights into the model’s limitations and identify potential avenues for further development. We now describe classes of errors that the models make. The three classes of mistakes most commonly identified are (1) missing or wrong predictions, (2) presence of non-comestibles, (3) non-descriptive predictions.
Missing descriptions
We noticed that this type of error occurs primarily in the T5 baseline. More specifically, the model will fail in predicting the masked tokens for longer recipes. This is a type of error that is frequent in the predictions in all settings and baselines. Some examples of what qualifies as wrong description are shown in Tab. 5. Given the instruction combine all ingredients except for the chicken, oil and cheese in a saucepan with ground truth input (tomato sauce; water; oregano; basil; thyme; garlic powder; salt; black pepper; bay leaf; lemon juice) and output seasoned tomato sauce, T5 predicts as input (chicken; oil; cheese) and output sauce. This is a clear example of wrong prediction where the model fails to "understand" that the ingredients explicitly mentioned in the instructions are not to be considered. Moreover, we can infer that in this case there is no internal understanding that the output should derive from the input, and we can deduct that the predicted sauce is identified from the instructions.
Presence of non comestibles
The last row of Table 5 is a clear example of presence of non comestibles in the predictions. Both GPT-3.5 + FT and T5+FT predict as input (chicken; fork), and consequently chicken with fork as output.
Non-descriptive predictions
Finally, GPT-3.5 predicts ingredients and ingredients in the bowl as input and output when given the instruction place everything in the bowl of an electric mixer with a dough hook. Although not completely wrong, the prediction is not informative enough to convey the nature of the intermediate food.
It is worth noting how BertScore is consistently higher in value than all other metrics. We explain this by the the fact that the predicted input-output pairs and the annotations are very close semantically. In this sense, this score represents an upper bound of our evaluation. On the other hand, EMA being a measure of precision of comparison between two strings is a lower bound.
8 Conclusion
In this work we propose PizzaCommonSense, a dataset for reasoning about intermediate input and output for each instruction step in cooking recipes. We argue that recipes often involve transformations that are not explicitly stated in the instructions, requiring the model to make inferences based on its understanding of the cooking process and the relationships between ingredients and actions. We provide a series of baselines to evaluate LLMs ability to perform this task these experiments demonstrate how challenging this task is. The results suggest that LLMs are capable of performing some basic reasoning tasks, but they still struggle with more complex tasks that require commonsense knowledge. By addressing these challenges, we can develop more powerful and versatile language models that can be used to reason about procedural texts. This could have a range of benefits, including improving the accuracy of automated recipe generation. The findings of this study have the potential to be generalized to a wide range of domains that rely on procedural texts.
9 Limitations
10 Ethical considerations
Data Collection
We performed the data collection using Amazon Mechanical Turk. We made sure annotators were fairly compensated by computing an average hourly wage above the US minimum wage. We maintain the anonymity of all data, ensuring that no personally identifiable information is captured from crowd workers. We rigorously curate the tasks and prompts used for data collection, meticulously avoiding any controversial or sensitive topics. This approach minimizes the potential for harm or misuse of the dataset.
Models and data source
The generative models are based on pre-trained language models, which may generate offensive content if prompted with inappropriate inputs. T5 and Flan-T5 are available on HuggingFace  [^1] https://huggingface.co/docs/transformers/. GPT-3.5  [^2] https://platform.openai.com/docs/models/gpt-3-5 was interacted with from the OpenAi API. our dataset was built from Recipe1M  [^3] http://pic2recipe.csail.mit.edu/ which is publicly available.
Appendix A Example of recipe
@@
Appendix B Example of predictions by T5
@@
Appendix C Example of predictions by Flan-T5
@@
Appendix D Example of predictions by GPT-3.5
@@
Appendix E Example of predictions by GPT-3.5 with demonstrations
@@
Appendix F Example of predictions by GPT-3.5 with fine-tuning

### v2

@@ 8 Related Work
Large datasets in the cooking domain, such as Food-101 Bossard et al. (2014) and Recipe1M Salvador et al. (2017b), have recently lead to significant advancements in food understanding. While these datasets are commonly used as benchmarks in computer vision (e.g., see Pham et al. (2021)), text-based natural language processing studies have explored areas like recipe text understanding via flow-graphs Mori et al. (2014a, b) and recipe parsing Chang et al. (2018); Jermsurawong and Habash (2015); Kiddon et al. (2015). We introduce a dataset that provides detailed input-output pairs and state changes in the cooking process in natural language, through step-by-step annotations. This detailed information is designed to develop models that can reason and articulate their reasoning paths at fine-grained level, an important feature for generating safe and coherent recipes.
@@
allows the model to focus on specific aspects of the task and to gradually build an understanding of the overall problem. One such prompting approach is the chain of thought prompting Wei et al. (2022), which prompts the language model to generate a series of intermediate steps that improve the reasoning capabilities in LLMs. Wang et al. took another step forward and sampled multiple reasoning paths and selected the most relevant output using majority voting. Kojima et al. further improved the reasoning of LLMs in a zero-shot manner by appending “Let’s think step by step” to the prompt. In contrast, our work explicitly asks the model to reason by trying to solve the sub-question at a fine-grained level. Most similar to our work is the work of Zhou et al. which decomposes questions into sub-questions and asks the language model to solve each sub-question sequentially.
@@
The reasoning process involves decomposing a multi-step reasoning task is inherently structured. While conventional natural language texts are generated in a 1-dimensional sequential order, the table has a 2-dimensional structure, which allows to reason horizontally and vertically at the same time. We argue that these features justify the choice of tabular data for the proposed task. To use an LLM for tabular data, the table must be serialized into a natural text representation. Proposed serialization formats include the simple list or sentence serializations Narayan et al. (2022); Borisov et al. (2022). Yin et al. (2020) also included the column data type in the serialized string. We used the serialization method by Wu et al. (2022).
9 Conclusion
In this work, we introduce PizzaCommonSense, a dataset for evaluating models’ understanding of cooking instructions, focusing on implicit ingredient transformations. We set baselines and demonstrate the challenging nature of the task through evaluations of LLMs. Our experiments underscore the significant limitations of LLMs when applied to this dataset, highlighting their inability to effectively handle complex reasoning that necessitates commonsense knowledge. These findings emphasize the need for advancements in model architecture and training methodologies to address these challenges. Overcoming these hurdles will enable the development of improved models capable of better understanding commonsense reasoning with procedural texts, with potential applications including autonomous agents using procedural information in scientific articles, industrial processes, DIY, as well as cooking.
Limitations
Ethical considerations
Data Collection
We performed the data collection using Amazon Mechanical Turk and data evaluation on CloudConnect Research. We made sure annotators were fairly compensated by calculating an average hourly wage above the US minimum wage. We maintain the anonymity of all data, ensuring that no personally identifiable information is captured from crowd workers. We rigorously curate the tasks and prompts used for data collection, meticulously avoiding any controversial or sensitive topics. This approach minimizes the potential for harm or misuse of the dataset.
Generative models
The generative models are based on pre-trained language models, which may generate offensive content if prompted with inappropriate inputs.
Acknowledgements
This research was supported by the Leverhulme Trust grant for the project ’Repurposing of Resources: from Everyday Problem Solving through to Crisis Management’ (RPG-2021-182). We also thank all reviewers for their insightful feedback.
Appendix A Dataset and Evaluation
Appendix B Models and data source
T5 and Flan-T5 are available on HuggingFace  [^1] https://huggingface.co/docs/transformers/. GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 and GPT-4333https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 were accessed from the OpenAi API. our dataset was built from Recipe1M  [^4] http://pic2recipe.csail.mit.edu/ which is publicly available.
Appendix C Additional tables
Table 6 represents the summary of PizzaCommonsense statistics.
@@ Appendix C Additional tables
Table 7 shows the (complete) quantitative evaluation for GPT-3.5+demo, GPT-3.5+FT and GPT-4+CoT.
Table 7: Quantitative evaluation for GPT-3.5+demo, GPT-3.5+FT and GPT-4+CoT. EMA is exact matching accuracy, B is Bleu, R is RougeL, M is Meteor and BS stands for BertScore. Higher is better.
@@
GPT3.5 + demo 22.3 36.9 32.5 9.20 20.0 87.2
GPT3.5 + FT 32.6 55.9 53.6 15.8 46.2 90.6
GPT-4 + CoT 26.7 51.4 50.9 9.14 40.7 88.9
Human 48.6 85.5 58.2 86.3 77.1 97.1
Appendix D Overlap Quantification
There is some overlap between individual cooking instructions in 3 sets. This is inevitable given that they are all pizza recipes which typically involves a limited set of steps: creating the base, applying sauce, adding toppings, and baking/chilling/setting. This is added to the fact that the task consist of describing “what we had before the action” and “what we get after the action”. Under this setting, our goal is to investigate if models can correctly reason even if the action is the same but the initial set of ingredients. We argue that it is the similar rationale of arithmetic reasoning datasets. The objective is to study compositional aspects of recipe instructions. Each stage (base, sauce, toppings) acts as a distinct "building block," allowing for easier segmentation and analysis of the construction and reasoning process.
The percentage overlap between training and validation sets for instructions is 5.47%, and 9.21% between training and testing. This overlap translates to ingredient lists with 3.91% and 6.26% overlap for training-validation and training-test sets, respectively. Similarly, the overlap in final outputs (resulting dishes) is 4.12% and 6.13% for training-validation and training-test sets. These duplicate instructions are mainly of the type "preheat the over at x" or "serve immediately" (some of the most common). In most cases, the instruction of the first type have as input "NA".
Appendix E Example of recipe
@@
Appendix F Example of predictions by T5
@@
Appendix G Example of predictions by Flan-T5
@@
Appendix H Example of predictions by GPT-3.5
@@
Appendix I Example of predictions by GPT-3.5 with demonstrations
@@
Appendix J Example of predictions by GPT-3.5 with fine-tuning
@@
Appendix K Example of predictions by GPT-4
<preheat oven to 350 degrees> <s> <in0> <s> preheat <s> NA
<mix broccoli, cheeses and seasonings> <s> <in1> <s> mix <s> cheese mixture
<spread 1 teaspoons of the cheese mixture onto each of the 48 crackers> <s> <in2> <s> spread <s> cheese-topped crackers
<place on baking sheet> <s> <in3> <s> place <s> crackers on baking sheet
<bake 5 minutes or until hot and bubbly> <s> <in4> <s> bake <s> baked cheese-topped crackers
<serve warm> <s> <in5> <s> serve <s> served cheese-topped crackers <n>
Appendix L Example of predictions by GPT-4+CoT
instructions <s> input <s> action <s> output <n>
preheat oven to 350 degrees <s> oven <s> preheat <s> preheated oven
mix broccoli , cheeses and seasonings <s> broccoli, cheeses, seasonings <s> mix <s> broccoli cheese mix
spread 1 teaspoons of the cheese mixture onto each of the 48 crackers <s> cheese mixture, crackers <s> spread <s> cheese-topped crackers
place on baking sheet <s> cheese-topped crackers <s> place <s> cheese-topped crackers on baking sheet bake 5 minutes or until hot and bubbly <s> cheese-topped crackers on baking sheet <s> bake <s> baked cheese-topped crackers
serve warm <s> baked cheese-topped crackers <s> serve <s> served baked cheese-topped crackers
Figure 3: Data collection interface on AMT.